OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? Answer 288 112 432 360

OpenStudy (anonymous):

360

OpenStudy (anonymous):

x = 0.8x + 72 x - 0.8x = 72 0.2x = 72 x = 72/0.2 x = 360